OpenStudy (anonymous):

together, lauren and aaron sold 456 raffle tickets. Lauren sold 24 more than aaron. How many did lauren sell?

OpenStudy (asnaseer):

lets say aaron sold 'x' tickets. then we know that lauren sold 'x+24'. total number of tickets sold is 456, so we can say: tickets sold by aaron + tickets sold by lauren = 456 x + 24 + x = 456

OpenStudy (asnaseer):

you should be able to solve from there...

OpenStudy (anonymous):

456 tickets / 2 people= 228 24 ticket difference / 2 people = 12 Lauren sold 228+12 = 240 Aaron sold 228-12 = 216

OpenStudy (asnaseer):

recall your previous problem, so we can simplify this to: 2x + 24 = 456 now take 24 away from both sides

OpenStudy (asnaseer):

can you do that step?

OpenStudy (anonymous):

2x+24-24=456-24 right?

OpenStudy (asnaseer):

correct, and this simplifies to: 2x = 432 so now divide both sides by 2...

OpenStudy (anonymous):

1= 216?!

OpenStudy (asnaseer):

x = 216 not 1 = 216 - nut I think you get the idea

OpenStudy (asnaseer):

*but

OpenStudy (anonymous):

then 216 + 24 right?!!

OpenStudy (asnaseer):

so now we know lauren sold x+24 = 216 +24 (as you said) - well done!

OpenStudy (anonymous):

yay! :)